# Temperature is All You Need for Generalization in Langevin Dynamics and other Markov Processes

### v1

@@ A.1 Preliminaries
[…]
@@ A.2 General Lemmas: Data processing inequality and generalized second laws of thermodynamics
[…]
@@
[…]
@@
[…]
@@
Therefore according to Lemma A.1,
@@
[…]
@@
Note we can assume $D_{\infty}\left(p_{n}\,\|\,p_{\infty}\right)<\infty$, since otherwise the claim holds trivially. Assume $\forall x:\left|\ln\frac{\mathrm{d}p_{n}}{\mathrm{d}p_{\infty}}(x)\right|\leq C$. Then for any $y$:
$\frac{\mathrm{d}p_{n+1}}{\mathrm{d}p_{\infty}}(y)$ $=\int\frac{\mathrm{d}p_{n}}{\mathrm{d}p_{\infty}}(x)P(y\mid x)\,\mathrm{d}p_{\infty}(x)$
$\leq e^{C}\int P(y\mid x)\,\mathrm{d}p_{\infty}(x)=e^{C}.$
@@
Let $0<t$ and let $\Delta t>0$ such that $t\in\Delta t\cdot\mathbb{N}$. Define $Y_{n}=X_{n\Delta t}$, then $Y_{n}$ is a discrete time Markov chain with marginals $p_{n\cdot\Delta t}$ and stationary distribution $p_{\infty}$, so Lemma A.2 and Lemma A.3 imply the results. ∎
Appendix B Proof of Theorem 2.7
In this section, we present the proof of Theorem 2.7, and the claims leading to it. We start by presenting the PAC-Bayes guarantees we rely on:
Theorem B.1 (Theorem 2.1 in [1], modified).
For any $\delta\in\left(0,1\right)$, for any data independent prior $\rho$,
$\mathbb{P}_{S\sim D^{N}}\left(\forall_{\hat{\rho}}\;\mathbb{E}_{\hat{\rho}}\left[E_{D}\left(h\right)-E_{S}\left(h\right)\right]\leq\sqrt{\frac{\mathrm{KL}\left(\hat{\rho}\,\middle\|\,\rho\right)+\ln\frac{1}{\delta}}{2N}}\right)\geq 1-\delta\,,$
where $\forall_{\hat{\rho}}$ is over all possible posterior distributions over predictors.
Theorem B.2 (Theorem 2.7 in [1], modified).
For any $\delta\in\left(0,1\right)$, for any data independent prior $\rho$, and any learning rule specified by a conditional probability $h|S\sim\hat{\rho}_{S}$,
$\mathbb{P}_{S\sim D^{N},h\sim\hat{\rho}_{S}}\left(E_{D}\left(h\right)-E_{S}\left(h\right)\leq\sqrt{\frac{D_{\infty}\left(\hat{\rho}_{S}\,\|\,\rho\right)+\ln\frac{1}{\delta}}{2N}}\right)\geq 1-\delta\,.$ (13)
The original theorem statements of Theorems B.1 and B.2 both include a variable $\lambda$ that we optimize over to arrive at these versions. For the single sample bound, we can do it after using $D_{\infty}\left(\hat{\rho}_{S}\,\|\,\rho\right)$ to uniformly bound $\ln\left(\mathop{}\!\mathrm{d}\hat{\rho}/\mathop{}\!\mathrm{d}\rho\right)$.
Remark B.3.
We use Theorems B.1 and B.2 to derive bounds in probability over the sample $S$. For bounds in expectation, one can similarly use Alquier et al. [1]’s Theorem 2.8.
In light of these theorems, all that remains is to bound the divegence terms.
  • Recall Claim 2.3.If $p,q,\mu,\nu$ are probability measures, and $p$ is Gibbs w.r.t $q$ with potential $\Psi<\infty$, then
@@
Using the Chain Rule and Claim 2.4, we derive the bounds of (1) and (2), as re-stated and established in the following lemma.
@@
where in the first inequality we used Claim 2.4. Similarly,
$D_{\infty}\left(p_{t}\,\|\,\nu\right)$ $=\operatorname{ess\,sup}_{p_{t}}\ln\frac{\mathop{}\!\mathrm{d}p_{t}}{\mathop{}\!\mathrm{d}\nu}=\operatorname{ess\,sup}_{p_{t}}\ln\frac{\mathop{}\!\mathrm{d}p_{t}}{\mathop{}\!\mathrm{d}p_{\infty}}\frac{\mathop{}\!\mathrm{d}p_{\infty}}{\mathop{}\!\mathrm{d}\nu}=D_{\infty}\left(p_{t}\,\|\,p_{\infty}\right)+D_{\infty}^{p_{t}}\left(p_{\infty}\,\|\,\nu\right)$
@@
Corollary 2.5 now follows from plugging in Claim 2.3 into Lemma B.4. And then Theorem 2.7 follows by plugging Corollary 2.5 into Theorems B.1 and B.2.
Instead of using the simple PAC-Bayes bounds of Theorems B.1 and B.2, we can just as well plug the divergence bound of 2.5 into tighter PAC-Bayes bounds to obtain tighter guarantees. For example, we could use:
@@
$\mathbb{P}_{S\sim D^{N}}\left(\forall_{\hat{\rho}\;}\mathrm{kl}\left(\mathbb{E}_{\hat{\rho}}E_{S}\left(h\right)\,\middle\|\,\mathbb{E}_{\hat{\rho}}E_{D}\left(h\right)\right)\leq\frac{\mathrm{KL}\left(\hat{\rho}\,\middle\|\,\rho\right)+\ln\frac{2\sqrt{N}}{\delta}}{N}\right)\geq 1-\delta\,,$
where $\mathrm{kl}\left(a\,\middle\|\,b\right)=a\ln\tfrac{a}{b}+(1-a)\ln\tfrac{1-a}{1-b}$ for $0\leq a,b\leq 1$ is the KL divergence for a Bernoulli random variable.
Or the high-probability variant:
@@
For any $\delta\in(0,1)$ and $N\geq 8$, for any data independent prior $\rho$, and any learning rule specified by a conditional probability $h|S\sim\hat{\rho}_{S}$,
$\mathbb{P}_{S\sim D^{N},h\sim\hat{\rho}_{S}}\left(\mathrm{kl}\left(E_{S}\left(h\right)\,\middle\|\,E_{D}\left(h\right)\right)\leq\frac{\ln\frac{\mathop{}\!\mathrm{d}\hat{\rho}_{S}}{\mathop{}\!\mathrm{d}\rho}(h)+\ln\frac{2\sqrt{N}}{\delta}}{N}\right)\geq 1-\delta\,.$
@@
Following and modifying the proof of Theorem 5 of Maurer [35], we start with the inequality $\mathbb{E}_{S}\left[e^{N\mathrm{kl}\left(E_{S}\left(h\right)\,\middle\|\,E_{D}\left(h\right)\right)}\right]\leq 2\sqrt{N}$ [35, Theorem 1], which holds for any $h$, and so also in expectation over $h$ w.r.t. $\rho$:
@@
Consider any distribution $D$ over $\mathcal{Z}$, function $f:\mathcal{H}\times\mathcal{Z}\to[0,1]$, and sample size $N>8$, any distribution $\nu$ over $\mathcal{H}$, and any discrete or continuous time process $\{h_{t}\in\mathcal{H}\}_{t\geq 0}$ (i.e. $t\in\mathbb{Z}_{+}$ or $t\in\mathbb{R}_{+}$) that is time-invariant Markov conditioned on $S$. Denote $p_{0}(\cdot;S)$ the initial distribution of the Markov process (that may depend on $S$). Let $p_{\infty}(\cdot;S)$ be any stationary distribution of the process conditioned on $S$, and $\Psi_{S}(h)\geq 0$ a non-negative potential function that can depend arbitrarily on $S$, such that $p_{\infty}(\cdot;S)$ is Gibbs w.r.t. $\nu$ with potential $\Psi_{S}$. Then:
@@
    $\mathrm{kl}\left(\mathbb{E}\left[E_{S}(h_{t})\middle|S\right]\,\middle\|\,\mathbb{E}\left[E_{D}(h_{t})\middle|S\right]\right)$ $\leq\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{N}}{{\delta}}}{N}$ (17)
@@
    $\mathbb{E}\left[E_{D}(h_{t})-E_{S}(h_{t})\middle|S\right]$ $\leq\sqrt{2\mathbb{E}\left[E_{S}\left(h_{t}\right)\mid S\right]\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{N}}{{\delta}}}{N}}$
    $\quad+2\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{N}}{{\delta}}}{N}$ (18)
@@
    $\mathrm{kl}\left(E_{S}(h_{t})\,\middle\|\,E_{D}(h_{t})\right)$ $\leq\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{N}}{{\delta}}}{N}$ (19)
    and so
    $E_{D}(h_{t})-E_{S}(h_{t})$ $\leq\sqrt{2E_{S}\left(h_{t}\right)\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{N}}{{\delta}}}{N}}$
    $\quad+2\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{N}}{{\delta}}}{N}$ (20)
@@
The inequalities (17) and (19) follow by plugging Corollary 2.5 into Theorems B.5 and B.6. For inequalities (18) and (20), we use the inequality $a\leq b+\sqrt{2b\mathrm{kl}\left(b\,\middle\|\,a\right)}+2\mathrm{kl}\left(b\,\middle\|\,a\right)$ for $a<b$. For (18), we use $a=E_{D}(h_{t})$ and $b=E_{S}(h_{t})$, which yields:
@@
$\leq E_{S}(h_{t})+\sqrt{2E_{S}(h_{t})\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{N}}{{\delta}}}{N}}$
$\quad+2\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{N}}{{\delta}}}{N},$
@@
Finally, we prove the equivalence statement made in Footnote 2:

### v2

@@ A.1 Preliminaries
We use bold lowercase letters (e.g. $\mathbf{x}\in\mathbb{R}^{d}$) to denote vectors, bold capital letters to denote matrices (e.g. $\mathbf{A}\in\mathbb{R}^{d\times d}$), and regular capital letters to denote random elements (e.g. $S,X,Y$). We may deviate from these conventions when it does not create confusion. Unless stated otherwise, all vectors are assumed to be column vectors. Specifically, we use $\mathbf{e}_{i}\in\mathbb{R}^{d}$, $i=1,\dots,d$, to denote the standard basis vector with $1$ in the $i^{th}$ entry, and 0 elsewhere. For a subset $\Omega\subseteq\mathbb{R}^{d}$, we denote by $\overline{\Omega}$, $\partial\Omega$, and $\Omega^{\circ}$, the closure, boundary, and interior of $\Omega$, respectively. In addition, we denote the volume of $B\subset\Omega$, when it is defined, by $\left\lvert{B}\right\rvert$}. With some abuse of notation, when $B$ is finite we denote its cardinality by $\left\lvert{B}\right\rvert$}. We use $\left\|\cdot\right\|$ for the standard Euclidean norm on $\mathbb{R}^{d}$. Then, the open Euclidean ball centered at $\mathbf{x}\in\mathbb{R}^{d}$ with radius $r>0$ is $B_{r}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\mathbb{R}^{d}\,\mid\,\left\|\mathbf{y}-\mathbf{x}\right\|<r\right\}$. In addition, we use $\mathbb{I}\left\{\cdot\right\}$ for the indicator function, and specifically for $A\subset\mathbb{R}^{d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, $\mathbb{I}_{A}\left\{x\right\}=\mathbb{I}\left\{\mathbf{x}\in A\right\}$. We denote the set of all probability measures over $\Omega$ by $\Delta\left(\Omega\right)$. For some $\mu\in\Delta\left(\Omega\right)$ with density $p$, with some abuse of notation we denote $p\in\Delta\left(\Omega\right)$, and $p\left(B\right)=\mu\left(B\right)$ for measurable $B\subseteq\Omega$. In addition, we use $\mathbb{E}_{X\sim p}$ or $\mathbb{E}_{p}$ to denote the expectation w.r.t $p$, and omit the subscript when it can be inferred. For two distributions $\mu,\nu$ with densities $p,q$ we denote by $\mathrm{KL}\left(\mu\,\middle\|\,\nu\right)=\mathrm{KL}\left(p\,\middle\|\,q\right)$ their KL-divergence (relative entropy). Furthermore, we use $H\left(\delta\right)=-\delta\ln\left(\delta\right)-\left(1-\delta\right)\ln\left(1-\delta\right)$, $\delta\in\left[0,1\right]$, for the binary entropy function (in nats). We denote the divergence of a vector field by $\nabla\cdot$, and the gradient and Laplacian of a scalar function by $\nabla$ and $\Delta=\nabla\cdot\nabla$, respectively. Given a domain $E\subset\mathbb{R}^{k}$ and $k\in\mathbb{Z}_{+}\cup\left\{\infty\right\}$, we denote by $\mathcal{C}^{k}\left(E\right)$ the set of real valued functions that are continuous over $\bar{E}$, and $k$-times continuously differentiable with continuous partial derivatives in $E$. In particular, $\mathcal{C}=\mathcal{C}^{0}$ is the set of continuous functions.
@@ A.2 General Lemmas: Data processing inequality and generalized second laws of thermodynamics
$\mathrm{KL}\left(p\,\middle\|\,q\right)$ $=\intop_{{\cal X}\times{\cal Y}}p\left(x,y\right)\ln\left(\frac{p\left(x,y\right)}{q\left(x,y\right)}\right)dxdy$
$=\intop_{{\cal X}\times{\cal Y}}p\left(x,y\right)\ln\left(\frac{p\left(y\mid x\right)p_{X}\left(x\right)}{q\left(y\mid x\right)q_{X}\left(x\right)}\right)dxdy$
$=\intop_{{\cal X}\times{\cal Y}}p\left(x,y\right)\ln\left(\frac{p_{X}\left(x\right)}{q_{X}\left(x\right)}\right)dxdy+\intop_{{\cal X}\times{\cal Y}}p\left(x,y\right)\ln\left(\frac{p\left(y\mid x\right)}{q\left(y\mid x\right)}\right)dxdy$
$=\intop_{{\cal X}\times{\cal Y}}p\left(y\mid x\right)p_{X}\left(x\right)\ln\left(\frac{p_{X}\left(x\right)}{q_{X}\left(x\right)}\right)dxdy+\intop_{{\cal X}\times{\cal Y}}p_{X}\left(x\right)p\left(y\mid x\right)\ln\left(\frac{p\left(y\mid x\right)}{q\left(y\mid x\right)}\right)dxdy$
$\left[\text{Fubini}\right]$ $=\intop_{{\cal X}}p_{X}\left(x\right)\ln\left(\frac{p_{X}\left(x\right)}{q_{X}\left(x\right)}\right)dx+\mathbb{E}_{X\sim p_{X}}\intop_{{\cal Y}}p\left(y\mid X\right)\ln\left(\frac{p\left(y\mid X\right)}{q\left(y\mid X\right)}\right)dy$
@@
$p_{n+1}\left(y\right)=\intop_{\Omega}P\left(y\mid x\right)p_{n}\left(x\right)dx\,.$
@@
$p_{\infty}\left(y\right)=\intop_{\Omega}P\left(y\mid x\right)p_{\infty}\left(x\right)dx\,.$
@@
Therefore according to Lemma˜A.1,
@@
$\mathrm{KL}\left(q\,\middle\|\,r\right)$ $=\intop_{\Omega\times\Omega}q\left(x,y\right)\ln\left(\frac{q\left(x,y\right)}{r\left(x,y\right)}\right)dxdy$
$=\intop_{\Omega\times\Omega}q\left(x,y\right)\ln\left(\frac{p_{n}\left(x\right)P\left(y\mid x\right)}{p_{\infty}\left(x\right)P\left(y\mid x\right)}\right)dxdy$
$=\intop_{\Omega\times\Omega}q\left(x,y\right)\ln\left(\frac{p_{n}\left(x\right)}{p_{\infty}\left(x\right)}\right)dxdy$
$=\intop_{\Omega\times\Omega}p_{n}\left(x\right)P\left(y\mid x\right)\ln\left(\frac{p_{n}\left(x\right)}{p_{\infty}\left(x\right)}\right)dxdy$
$\left[\text{Fubini}\right]$ $=\intop_{\Omega}p_{n}\left(x\right)\ln\left(\frac{p_{n}\left(x\right)}{p_{\infty}\left(x\right)}\right)dx$
@@
Let $p,q$ be some probability measures such that $\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}$ exists. By definition,
$D_{\infty}\left(p\,\|\,q\right)=\operatorname{ess\,sup}_{q}\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}=\inf\left\{c\in\mathbb{R}\,\mid\,q\left(\left\{x\,\mid\,\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}>c\right\}\right)=0\right\}\,.$
Let $C\in\mathbb{R}$ and suppose that for all measurable $A\subset\mathcal{X}$, $p\left(A\right)\leq e^{C}q\left(A\right)$. Assume by way of contradiction that $D_{\infty}\left(p\,\|\,q\right)>C$, that is, that there exists $c>C$ such that
$q\left(\left\{x\,\mid\,\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}>c\right\}\right)>0\,.$
Denote
$A=\left\{x\,\mid\,\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}>c\right\}\,$,
then
$p\left(A\right)=\intop_{A}\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}\mathop{}\!\mathrm{d}q>e^{c}q\left(A\right)>e^{C}q\left(A\right)\,$,
in contradiction to the assumption. Therefore, for all $C$ such that $p\left(A\right)\leq e^{C}q\left(A\right)$ for all measurable $A$, $C\geq D_{\infty}\left(p\,\|\,q\right)$. We can now show the claim.
Let $P(\mathop{}\!\mathrm{d}y\mid x)$ be the processes’ transition kernel (in measure form). We can assume $D_{\infty}\left(p_{n}\,\|\,p_{\infty}\right)<\infty$, since otherwise the claim holds trivially. Let $A$ be measurable, then by definition,
$p_{n+1}\left(A\right)$ $=\intop P\left(A\mid x\right)dp_{n}\left(x\right)=\intop P\left(A\mid x\right)\frac{\mathop{}\!\mathrm{d}p_{n}}{\mathop{}\!\mathrm{d}p_{\infty}}\left(x\right)\mathop{}\!\mathrm{d}p_{\infty}\left(x\right)$
$\leq e^{D_{\infty}\left(p_{n}\,\|\,p_{\infty}\right)}\intop P\left(A\mid x\right)\mathop{}\!\mathrm{d}p_{\infty}\left(x\right)=e^{D_{\infty}\left(p_{n}\,\|\,p_{\infty}\right)}p_{\infty}\left(A\right)\,$,
@@
Let $0<t$ and let $\Delta t>0$ such that $t\in\Delta t\cdot\mathbb{N}$. Define $Y_{n}=X_{n\Delta t}$, then $Y_{n}$ is a discrete time Markov chain with marginals $p_{n\cdot\Delta t}$ and stationary distribution $p_{\infty}$, so Lemma˜A.2 and Lemma˜A.3 imply the results. ∎
Appendix B Proof of Theorem˜2.7 and its Related Claims in Section˜2
In this section, we present the proof of Theorem˜2.7, the claims leading to it, and some of its generalizations.
B.1 Derivation of Corollary˜2.5
Recall Claim˜2.3.If $p,q,\mu,\nu$ are probability measures, and $p$ is Gibbs w.r.t $q$ with potential $\Psi<\infty$, then
@@
Using the Chain Rule and Claim˜2.4, we derive the bounds of (1) and (2), as re-stated and established in the following lemma.
@@
where in the first inequality we used Claim˜2.4. Similarly,
$D_{\infty}\left(p_{t}\,\|\,\nu\right)$ $=\operatorname{ess\,sup}_{p_{t}}\ln\frac{\mathop{}\!\mathrm{d}p_{t}}{\mathop{}\!\mathrm{d}\nu}=\operatorname{ess\,sup}_{p_{t}}\ln\frac{\mathop{}\!\mathrm{d}p_{t}}{\mathop{}\!\mathrm{d}p_{\infty}}\frac{\mathop{}\!\mathrm{d}p_{\infty}}{\mathop{}\!\mathrm{d}\nu}\leq D_{\infty}\left(p_{t}\,\|\,p_{\infty}\right)+D_{\infty}^{p_{t}}\left(p_{\infty}\,\|\,\nu\right)$
@@
Corollary˜2.5 now follows from plugging in Claim˜2.3 into Lemma˜B.1.
Given these bounds on the divergences, All that remains in order to prove Theorem˜2.7 is plugging Corollary˜2.5 into a PAC-Bayes bound.
B.2 In-Expectation PAC-Bayes Bounds
@@
$\mathbb{P}_{S\sim D^{N}}\left(\forall_{\hat{\rho}\;}\mathrm{kl}\left(\mathbb{E}_{h\sim\hat{\rho}}E_{S}\left(h\right)\,\middle\|\,\mathbb{E}_{h\sim\hat{\rho}}E_{D}\left(h\right)\right)\leq\frac{\mathrm{KL}\left(\hat{\rho}\,\middle\|\,\rho\right)+\ln\frac{2\sqrt{N}}{\delta}}{N}\right)\geq 1-\delta\,$,
where $\mathrm{kl}\left(a\,\middle\|\,b\right)=a\ln\tfrac{a}{b}+(1-a)\ln\tfrac{1-a}{1-b}$ for $0\leq a,b\leq 1$ is the KL divergence for a Bernoulli random variable, and $\hat{\rho}$ denotes a posterior distribution.
B.3 Single-Sample PAC-Bayes Bounds
Theorem˜B.2 can be viewed as a bound in expectation over the draw from the posterior, which corresponds to the traditional PAC-Bayes view of considering the expected error of a randomized predictor. But it is actually possible to get guarantees for a single draw from this predictor, which is more appropriate when we view the randomness as part of the training algorithm, that then outputs a single deterministic predictor (chosen at random). High probability guarantees for a single draw from the posterior were shown by Alquier et al. [1] based on Catoni [9] and also discussed by Dziugaite and Roy [16]. Here we present a tight version based on a simple modification to Maurer’s proof [43].
@@ B.3 Single-Sample PAC-Bayes Bounds
For any $\delta\in(0,1)$ and $N\geq 8$, for any data independent prior $\rho$, and any learning rule specified by a conditional probability $h|S\sim\hat{\rho}_{S}$ such that $\rho\ll\hat{\rho}_{S}$ $S$-a.s.,
$\mathbb{P}_{S\sim D^{N},h\sim\hat{\rho}_{S}}\left(\mathrm{kl}\left(E_{S}\left(h\right)\,\middle\|\,E_{D}\left(h\right)\right)\leq\frac{\ln\frac{\mathop{}\!\mathrm{d}\hat{\rho}_{S}}{\mathop{}\!\mathrm{d}\rho}(h)+\ln\frac{2\sqrt{N}}{\delta}}{N}\right)\geq 1-\delta\,$,
and so, by the definition of $D_{\infty}\left(\hat{\rho}_{S}\,\|\,\rho\right)$,
$\mathbb{P}_{S\sim D^{N},h\sim\hat{\rho}_{S}}\left(\mathrm{kl}\left(E_{S}\left(h\right)\,\middle\|\,E_{D}\left(h\right)\right)\leq\frac{D_{\infty}\left(\hat{\rho}_{S}\,\|\,\rho\right)+\ln\frac{2\sqrt{N}}{\delta}}{N}\right)\geq 1-\delta\,.$
@@
Following and modifying the proof of Theorem 5 of Maurer [43], we start with the inequality $\mathbb{E}_{S}\left[e^{N\mathrm{kl}\left(E_{S}\left(h\right)\,\middle\|\,E_{D}\left(h\right)\right)}\right]\leq 2\sqrt{N}$ [43, Theorem 1], which holds for any $h$, and so also in expectation over $h$ w.r.t. $\rho$:
@@
Consider any distribution $D$ over $\mathcal{Z}$, function $f:\mathcal{H}\times\mathcal{Z}\to[0,1]$, and sample size $N\geq 8$, any distribution $\nu$ over $\mathcal{H}$, and any discrete or continuous time process $\{h_{t}\in\mathcal{H}\}_{t\geq 0}$ (i.e. $t\in\mathbb{Z}_{+}$ or $t\in\mathbb{R}_{+}$) that is time-invariant Markov conditioned on $S$. Denote $p_{0}(\cdot;S)$ the initial distribution of the Markov process (that may depend on $S$). Let $p_{\infty}(\cdot;S)$ be any stationary distribution of the process conditioned on $S$, and $\Psi_{S}(h)\geq 0$ a non-negative potential function that can depend arbitrarily on $S$, such that $p_{\infty}(\cdot;S)$ is Gibbs w.r.t. $\nu$ with potential $\Psi_{S}$. Then:
@@
    $\mathrm{kl}\left(\mathbb{E}\left[E_{S}(h_{t})\middle|S\right]\,\middle\|\,\mathbb{E}\left[E_{D}(h_{t})\middle|S\right]\right)$ $\leq\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}$ (16)
@@
    $\mathbb{E}\left[E_{D}(h_{t})-E_{S}(h_{t})\middle|S\right]$ $\leq\sqrt{2\mathbb{E}\left[E_{S}\left(h_{t}\right)\mid S\right]\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}}$
    $\quad+2\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}$ (17)
@@
    $\mathrm{kl}\left(E_{S}(h_{t})\,\middle\|\,E_{D}(h_{t})\right)$ $\leq\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}$ (18)
    and so, when $E_{S}(h_{t})<E_{D}(h_{t})$
    $E_{D}(h_{t})-E_{S}(h_{t})$ $\leq\sqrt{2E_{S}\left(h_{t}\right)\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}}$
    $\quad+2\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}$ (19)
Lemma B.5.
Let $a,b\in\left[0,1\right]$. Then
$b\leq a+\sqrt{2a\mathrm{kl}\left(a\,\middle\|\,b\right)}+2\mathrm{kl}\left(a\,\middle\|\,b\right)\,.$ (20)
@@
The KL divergence is non-negative, so it suffices to consider the case that $b\geq a$. Defining $\varphi:\left[0,1-a\right]\to\mathbb{R}$ as
$\varphi\left(u\right)=\frac{u^{2}}{2\left(a+u\right)}\,$,
it can be readily checked by differentiation that for all $u\in\left[0,1-a\right]$,
$\mathrm{kl}\left(a\,\middle\|\,a+u\right)\geq\varphi\left(u\right)\,.$
In particular, for $u=b-a\in\left[0,1-a\right]$,
$\mathrm{kl}\left(a\,\middle\|\,b\right)\geq\frac{\left(b-a\right)^{2}}{2b}\,.$ (21)
Next, we consider the following inequality
$2u^{2}+\sqrt{2a}u+a-b\geq 0\,,\;u\geq 0\,.$ (22)
Solving for $u$, it turns out that the inequality holds when
$u\geq\frac{\sqrt{8b-6a}-\sqrt{2a}}{4}\,.$ (23)
In addition, under the assumption that $b\geq a$,
$\frac{\sqrt{8b-6a}-\sqrt{2a}}{4}\leq\sqrt{\frac{\left(b-a\right)^{2}}{2b}}\,.$ (24)
Combining (21), (23), and (24), $u=\sqrt{\mathrm{kl}\left(a\,\middle\|\,b\right)}$ solves (22) implying (20). ∎
Proof.
The inequalities (16) and (18) follow by plugging Corollary˜2.5 into Theorems˜B.2 and B.3. For inequalities (17) and (19), we use (20). For (17), we use $a=E_{D}(h_{t})$ and $b=E_{S}(h_{t})$, which yields:
@@
$\leq E_{S}(h_{t})+\sqrt{2E_{S}(h_{t})\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N}}$
$\quad+2\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}_{p_{0}}\Psi_{S}(h_{0})+\ln\nicefrac{{2\sqrt{N}}}{{\delta}}}{N},$
@@
Remark B.7.
In order to get the version in Theorem˜2.7 we use the upper bound of Pinsker’s inequality, i.e. that for all $a,b\in\left(0,1\right)$
${\left\lvert{a-b}\right\rvert}\leq\sqrt{\frac{1}{2}\mathrm{kl}\left(a\,\middle\|\,b\right)}\,$,
and simplify $\ln\frac{2\sqrt{N}}{\delta}\leq\ln\frac{N}{\delta}$ as $N\geq 8$.
Finally, we prove the equivalence statement made in Footnote˜6: